OpenStudy (anonymous):

how do i type all real numbers except 0 and -2 in interval notation

zepdrix (zepdrix):

Intervals? Ummm :) So we're going to have the `union` of a few sets.

zepdrix (zepdrix):

Woops, you listed them 0 before -2, my bad. lemme fix that.

zepdrix (zepdrix):

We go from left to right, so we'll start from negative infinity until we get to -2,\[\Large\rm (-\infty,-2)\]We put a rounded bracket around the -2 to show that we're excluding that value, and note that we ALWAYS put rounded around infinity.

zepdrix (zepdrix):

Then we want the values from -2 to 0, excluding -2 and 0.\[\Large\rm (-2,0)\]

zepdrix (zepdrix):

Andddddd also all the values from 0 up to infinity, excluding the 0.\[\Large\rm (0,\infty)\]And then we just take the union of those three sets! :)

OpenStudy (anonymous):

what is the union?

zepdrix (zepdrix):

The union of the first two will be:\[\Large\rm (-\infty,-2) \cup (-2,0)\]

zepdrix (zepdrix):

And we also want to include the third piece.\[\Large\rm (-\infty,-2) \cup (-2,0) \cup (0,\infty)\]Soooo.. yah...